What will be energy of proton if it is moving with 97% speed of light?
1
Expert's answer
2020-06-04T10:05:24-0400

Put that "\alpha=0.97=v/c." Then the energy of such a photon will be


"E=\sqrt{p^2c^2+(m_0c^2)^2}=\sqrt{\frac{m_0^2\alpha^2c^4}{1-\alpha^2}+(m_0c^2)^2}=\\\space\\
=\frac{m_0c^2}{\sqrt{1-\alpha^2}}=\frac{1.673\cdot10^{-27}(3\cdot10^8)^2}{\sqrt{1-0.97^2}}=1.671\cdot10^{-17}\text{ J},"

or 104.3 eV.